OpenStudy (anonymous):

Just looking to make sure this problem was done correctly. Factor the following polynomial completely: (3x-5)(x-3)-(x-4)(x-3) The answer I got was 2x-26

OpenStudy (anonymous):

I wanted to make sure the answer was correct, so if you wanted to double check it that would be awesome.

OpenStudy (anonymous):

Yeah, the foil method is what i did.

OpenStudy (btaylor):

well, there is a gcf you can factor out without having to multiply it all and then re-factor. Do you see the common factor?

OpenStudy (anonymous):

x

OpenStudy (btaylor):

if you pull out a (x-3), you get (x-3)[(3x-5)-(x-4)] (x-3)(3x-x-5+4) (x-3)(2x-1)

OpenStudy (anonymous):

Here is what i did:\[(3x-5)(x-3)=3x ^{2}-9x-41 (x-4)(x-3)=x ^{2}-7x+12\]

OpenStudy (anonymous):

Then subtracted the trinomials.

OpenStudy (anonymous):

I know how the foil method works.

OpenStudy (btaylor):

@StefaniTriscuit I found your error. -5x - 9x = -14x , not -9. Then, -5 x -3 = 15 , not -41. If you fix that, then you've got it right.

OpenStudy (anonymous):

Ah! haha, thank you! so obvious now!

OpenStudy (btaylor):

you're welcome!